OpenStudy (anonymous):

PLEASE HELP WILL GIVE A MEDAL Find the standard form of the equation of the parabola with a focus at (0, -10) and a directrix at y = 10

OpenStudy (jdoe0001):

 notice that focus point and that directrix any ideas on where the vertext is at?

OpenStudy (anonymous):

0,0? i really don't know

OpenStudy (jdoe0001):

vertex is half-way between the directrix and focus point

OpenStudy (jdoe0001):

but you'd be correct in this case, half-way between those 2 fellows, is 0,0 keep in mind that the parabola opens towards the focus point in this case the focus point is down below, so the parabola opens dowwards notice the distance from the vertex to either the focus or directrix? that'd the distance "p" the parabola goes down, thus "p" will be negative and the equation will have a squared "x" variable thus \(\large {\begin{array}{llll} (x-{\color{brown}{ h}})^2=4p(y-{\color{blue}{ k}})\\ \end{array} \qquad \begin{array}{llll} vertex\ ({\color{brown}{ h}},{\color{blue}{ k}})\\ p=\textit{distance from vertex to }\\ \qquad \textit{ focus or directrix} \end{array} }\) so, now you know what (h,k) are, and "p" as well, thus plug them in :)